# Charge radii measurements of exotic tin isotopes in the proximity of and

### v1

@@ Acknowledgements.
We thank the ISOLDE technical group for their professional assistance. Useful comments from Ante Ravlić are gratefully acknowledged. This work was supported by ERC Consolidator Grant No. 648381 (FNPMLS); STFC grants ST/L005794/1, ST/L005786/1, (ST/L005670/1, ST/P004423/1, and Ernest Rutherford Grant No. ST/L002868/1; Projects GOA 15/010 and C14/22/104 from KU Leuven; the FWO-Vlaanderen project G080022N (Belgium); the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under grants DE-SC0021176 and DE-SC0013365, DE-SC0018083 and DE-SC0023175 (Office of Advanced Scientific Computing Research and Office of Nuclear Physics, Scientific Discovery through Advanced Computing); the Max-Planck Society; the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 279384907 – SFB 1245; the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 101020842); the German Federal Ministry for Education and Research under Contract No. 05P15RDCIA and 05P24RD4; the Helmholtz International Center for FAIR (HIC for FAIR) within the LOEWE program by the State of Hesse; the European Union Grant Agreement 654002 and 262010 (ENSAR); National Key R&D Program of China (Contract No: 2018YFA0404403); the National Natural Science Foundation of China (No:11875073); NSERC under grants SAPIN-2018-00027, RGPAS-2018-522453, SAPIN-2024-0003, and the Arthur B. McDonald Canadian Astroparticle Physics Research Institute. TRIUMF receives funding via a contribution through the Digital Research Alliance of Canada. Computations were performed with an allocation of computing resources on Cedar at WestGrid and Compute Canada. JK acknowledges support from a Feodor-Lynen postdoctoral research fellowship funded by the Alexander-von-Humboldt Foundation. We acknowledge the computing resources provided by MIT and the Texas A&M high-performance computing cluster. TO, NS, and YT acknowledge the support by “Program for Promoting Researches on the Supercomputer Fugaku” (JPMXP1020200105, hp220174, hp210165). TM and NS acknowledge the support by JST ERATO Grant No. JPMJER2304, Japan. PGR thanks the regional computing center of the university Erlangen (RRZE) for support. This work was performed under the auspices of the U.S. Department of Energy, Office of Nuclear Physics (DOE NP) by Oak Ridge National Laboratory under Contract No. DE-AC05-00OR22725. DLB is supported in part through grants from the Romanian Ministry of Research, Innovation and Digitalization under contracts No. PN 23 21 01 06 and ELI-RO-RDI-2024-007 (ELITE). The data analysis used NumPy [72], SciPy [73], Pandas [74, 75], and iMinuit [76, 77]. The figures were produced using Matplotlib [78].Appendix A Supplemental Material
VS-IMSRG calculations
@@
The VS-IMSRG calculations [32, 33] start from a Hamiltonian with two- and three-nucleon interactions represented in a space of 15 harmonic-oscillator shells with frequency MeV and the sufficiently large truncation for 3N matrix elements [79]. In the current study, we employed two chiral interactions, 1.8/2.0 (EM) [37, 38] and N2LOGO(394) [39]. Combined with the ensemble normal ordering method [80] to capture effects of 3N forces between valence nucleons, we decouple a valence space Hamiltonian using the Magnus expansion detailed in Ref. [81], where operators are truncated at the two-body level, the IMSRG(2) approximation. The radius operators are consistently transformed, and the charge radii are computed as in Ref. [82].
@@
In the current study, three different valence spaces were adopted: The one-major shell space (above a 80Zr core), the neutron orbits above a 100Sn core, and the neutron above a 114Sn core. The VS-IMSRG calculations were performed with the imsrg++ code [83], and the effective Hamiltonians for the 100Sn core space were diagonalized with KSHELL code [84]. Since the exact diagonalizations of the valence-space Hamiltonians within the -shell space are not tractable, these calculations required the use of the recently developed quasi-particle vacua for nuclear shell-model calculations QVSM [85].
Experimental methodology
@@
For the CRIS experiment, beams of stable and radioactive tin isotopes were produced from the bombardment of a heated LaCx target (1500-2000 ∘C) with a 1.4 GeV proton beam at the ISOLDE-CERN facility. The nuclear reaction recoils escape the target material through diffusion and effusion and enter a heated transfer line to an ion source. There, the tin atoms were selectively ionized with the resonant ionization laser ion source (RILIS) [86], and electrostatically accelerated to 40 keV. The ion-beam is steered through the high-resolution mass separator (HRS) magnets for isotope separation (). The separated beam is then sent into a cooler buncher linear Paul trap (ISCOOL) [87]. The ion bunches are released with a bunch length of 5 , and transported to the CRIS beamline, described in detail elsewhere [25]. Subsequently the ions were neutralized with sodium vapour at 230-250 ∘C within a charge-exchange-cell (CEC). The fraction of neutral atoms within the beam varied between 10–30 %. Any remaining ions were removed by electrostatic deflectors after the CEC. The atoms entered the interaction region held at a high vacuum of mbar to reduce the probability of collisional ionization of the neutral atom beam. Within the interaction region, each atom bunch was collinearly overlapped and synchronized in time with a series of laser pulses at 100 Hz repetition rate, which step-wise excited and non-resonantly ionized the atomic bunch. The two atomic transitions probed for studying the isotope shifts are (284 nm) and (281 nm). The two transitions were studied using the three- and two-step ionization schemes, as described in Ref. [34, 88]. These transitions were scanned with a speed below 5 MHz/s across the resonance peak using an injection-seeded single-mode titanium sapphire laser, operating at a repetition rate of 1 kHz, with a 20 MHz bandwidth in the fundamental output. The resonantly ionized ions are deflected onto an ETP DM291 MagneTOFTM detector with a timing resolution of 0.4 ns for event-per-event ion counting.
@@
Example spectra of the Doppler corrected resonances observed when probing the transition for the even-even () isotopes are presented in Fig. 3, revealing the isotopic shift which contains the information of the changes in the nuclear charge radii .
@@
Throughout the experiment, the ion beam acceleration voltage was measured, allowing accurate Doppler-correction of the spectra. The wavelength of the scanned step was regularly measured using a WSU-2 HighFinesse wavemeter, which simultaneously recorded a reference wavelength from a stabilized laser diode (Toptica DLC DL PRO 780) locked to the Doppler-free hyperfine transition of 87Rb (TEM CoSy) for correcting drifts of the wavemeter’s interferometer. Furthermore, data taking on the isotopes of interest was interleaved throughout the experiment campaign with measurements of selected reference isotope 112Sn to ensure any additional drifts were taken into account when determining the isotope shifts.
@@
The transition in tin was predominantly used throughout the experiment, allowing spectroscopy of 21 tin isotopes ranging from 104Sn to 124Sn. Complementary measurements were performed using the transition down to 106Sn. During the spectroscopy of 114,116Sn using the transition, the laser power of the first step was increased by a factor five for enhancing ionization efficiency while increasing the linewidth from 230 MHz to 350 MHz total FWHM. Reference measurements of both 112Sn and 120Sn were performed under these identical conditions. Resonances were fitted using a Lorentz profile with free linewidth, since no Gaussian contribution was observed, with an average linewidth of 230 MHz for all isotopes except those acquired using a higher laser power. As an example, the fitted spectra extracted using the transition of the even-even isotopes are presented in Fig. 3.
| CRIS | COLLAPS | |||||
| (MHz) | (MHz) | (MHz) | (MHz) | |||
| 104 | 0 | |||||
| 105 | 5/2+ | |||||
| 106 | 0 | |||||
| 107 | 5/2+ | |||||
| 108 | 0 | |||||
| 109 | 5/2+ | |||||
| 110 | 0 | |||||
| 111 | 7/2+ | |||||
| 112 | 0 | 0 | 0 | |||
| 113 | 1/2+ | |||||
| 114 | 0 | |||||
| 115 | 1/2+ | |||||
| 116 | 0 | |||||
| 117 | 1/2+ | |||||
| 118 | 0 | |||||
| 119 | 1/2+ | |||||
| 120 | 0 | |||||
| 121 | 3/2+ | |||||
| 122 | 0 | |||||
| 123 | 11/2- | |||||
| 124 | ] | |||||
| 125 | 11/2- | |||||
| 126 | 0 | ] | ||||
| 127 | 11/2- | |||||
| 128 | 0 | |||||
| 129 | 3/2+ | |||||
| 130 | ||||||
| 131 | 3/2+ | |||||
| 132 | ||||||
| 133 | 7/2- | |||||
| 134 | 0 | |||||
@@
The determined isotope shifts relative to 124Sn are provided in Tab. 2 for both transitions. In order to obtain from the isotope shifts measured at CRIS relative to 112Sn, the following procedure was used. First, the isotope shifts relative to 124Sn were calculated and the statistical uncertainties of the isotope of interest and that of the new reference 124Sn were added in quadrature. To determine the total uncertainty, we combined the statistical uncertainty with the systematic uncertainty specific to each isotope. In this particular case, the systematic uncertainty was not derived from a potential voltage offset. Instead, it stemmed from laser drifts that could occur during the measurement of different isotopes, whether in positive or negative directions. For the isotope 112Sn, we used the systematic uncertainty associated with 124Sn. This total uncertainty was then used to perform the King plot for the individual transitions. For the COLLAPS data, only the statistical uncertainty was taken into account, since the systematic uncertainty is dominated by the scaling uncertainty of the voltage divider that transfers into a mass-shift like behavior in the isotopic chain and is thus captured by the King-plot procedure. The final value for as listed in Table 1 is then obtained as the weighted mean of the measurements of the four transitions. The final uncertainty has been chosen as the smallest uncertainty that was obtained from one specific transition, but no further reduction was considered due to the strong correlations via the respective field shift uncertainty. It was found that the size of this uncertainty covers approximately the observed range of variation of the results in the four transitions.
@@
The mass shift and field shift factors obtained in the King plots, which are required to obtain the values in the respective transitions, are summarized in Table LABEL:table:atomic_factors. The differential mean-square charge radii and their uncertainties for each transition individually can be obtained using the values in Tab. 2 and Tab. LABEL:table:atomic_factors, according to with the mass factor . In the evaluation of the uncertainty on , the optimized ensures that no correlations between and exist in the error propagation.

### v2

@@ Acknowledgements.
We thank the ISOLDE technical group for their professional assistance. Useful comments from Ante Ravlić are gratefully acknowledged. This work was supported by ERC Consolidator Grant No. 648381 (FNPMLS); STFC grants ST/L005794/1, ST/L005786/1, (ST/L005670/1, ST/P004423/1, and Ernest Rutherford Grant No. ST/L002868/1; Projects GOA 15/010 and C14/22/104 from KU Leuven; the FWO-Vlaanderen project G080022N, G053221N, I002619N and I001323N (Belgium); the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under grants DE-SC0021176 and DE-SC0013365, DE-SC0018083 and DE-SC0023175 (Office of Advanced Scientific Computing Research and Office of Nuclear Physics, Scientific Discovery through Advanced Computing); the Max-Planck Society; the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 279384907 – SFB 1245; the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 101020842); the German Federal Ministry for Education and Research under Contract No. 05P15RDCIA and 05P24RD4; the Helmholtz International Center for FAIR (HIC for FAIR) within the LOEWE program by the State of Hesse; the European Union Grant Agreement 654002 and 262010 (ENSAR); National Key R&D Program of China (Contract No: 2018YFA0404403); the National Natural Science Foundation of China (No:11875073); NSERC under grants SAPIN-2018-00027, RGPAS-2018-522453, SAPIN-2024-0003, and the Arthur B. McDonald Canadian Astroparticle Physics Research Institute. TRIUMF receives funding via a contribution through the Digital Research Alliance of Canada. Computations were performed with an allocation of computing resources on Cedar at WestGrid and Compute Canada. JK acknowledges support from a Feodor-Lynen postdoctoral research fellowship funded by the Alexander-von-Humboldt Foundation. We acknowledge the computing resources provided by MIT and the Texas A&M high-performance computing cluster. TO, NS, and YT acknowledge the support by “Program for Promoting Researches on the Supercomputer Fugaku” (JPMXP1020200105, hp220174, hp210165). DTY acknowledges support from the Franco-Bulgarian Hubert Curien partnership Rila no. 51315QM and no. KP-06-RILA/4. TM and NS acknowledge the support by JST ERATO Grant No. JPMJER2304, Japan. PGR thanks the regional computing center of the university Erlangen (RRZE) for support. This work was performed under the auspices of the U.S. Department of Energy, Office of Nuclear Physics (DOE NP) by Oak Ridge National Laboratory under Contract No. DE-AC05-00OR22725. DLB is supported in part through grants from the Romanian Ministry of Research, Innovation and Digitalization under contracts No. PN 23 21 01 06 and ELI-RO-RDI-2024-007 (ELITE). D. T. Y. acknowledges support from the Franco-Bulgarian Hubert Curien partnership Rila No. 51315QM and No. KP-06-RILA/4.Appendix A Supplemental Material
A.1 VS-IMSRG calculations
@@
The VS-IMSRG calculations [Hergert:2015awm, Stro19ARNPS] start from a Hamiltonian with two- and three-nucleon interactions represented in a space of 15 harmonic-oscillator shells with frequency MeV and the sufficiently large truncation for 3N matrix elements [Miya22Heavy]. In the current study, we employed two chiral interactions, 1.8/2.0 (EM) [Entem2003, Hebe11fits] and N2LOGO(394) [Jia20]. Combined with the ensemble normal ordering method [Stro17ENO] to capture effects of 3N forces between valence nucleons, we decouple a valence space Hamiltonian using the Magnus expansion detailed in Ref. [Morris2015], where operators are truncated at the two-body level, the IMSRG(2) approximation. The radius operators are consistently transformed, and the charge radii are computed as in Ref. [Hagen2015].
@@
In the current study, three different valence spaces were adopted: The one-major shell space (above a 80Zr core), the neutron orbits above a 100Sn core, and the neutron above a 114Sn core. The VS-IMSRG calculations were performed with the imsrg++ code [Stroberg++], and the effective Hamiltonians for the 100Sn core space were diagonalized with KSHELL code [Shimizu2019]. Since the exact diagonalizations of the valence-space Hamiltonians within the -shell space are not tractable, these calculations required the use of the recently developed quasi-particle vacua for nuclear shell-model calculations QVSM [Shimizu2021].
A.2 Experimental methodology and the determination of nuclear charge radii
@@
For the CRIS experiment, beams of stable and radioactive tin isotopes were produced from the bombardment of a heated LaCx target (1500-2000 ∘C) with a 1.4 GeV proton beam at the ISOLDE-CERN facility. The nuclear reaction recoils escape the target material through diffusion and effusion and enter a heated transfer line to an ion source. There, the tin atoms were selectively ionized with the resonant ionization laser ion source (RILIS) [fedosseev2017ion], and electrostatically accelerated to 40 keV. The ion-beam is steered through the high-resolution mass separator (HRS) magnets for isotope separation (). The separated beam is then sent into a cooler buncher linear Paul trap (ISCOOL) [mane2009ion]. The ion bunches are released with a bunch length of 5 , and transported to the CRIS beamline, described in detail elsewhere [Kos20]. Subsequently the ions were neutralized with sodium vapour at 230-250 ∘C within a charge-exchange-cell (CEC). The fraction of neutral atoms within the beam varied between 10–30 %. Any remaining ions were removed by electrostatic deflectors after the CEC. The atoms entered the interaction region held at a high vacuum of mbar to reduce the probability of collisional ionization of the neutral atom beam. Within the interaction region, each atom bunch was collinearly overlapped and synchronized in time with a series of laser pulses at 100 Hz repetition rate, which step-wise excited and non-resonantly ionized the atomic bunch. The two atomic transitions probed for studying the isotope shifts are (284 nm) and (281 nm). The two transitions were studied using the three- and two-step ionization schemes, as described in Ref. [Fredrik2020, parnefjord2021evolution]. These transitions were scanned with a speed below 5 MHz/s across the resonance peak using an injection-seeded single-mode titanium sapphire laser, operating at a repetition rate of 1 kHz, with a 20 MHz bandwidth in the fundamental output. The resonantly ionized ions are deflected onto an ETP DM291 MagneTOFTM detector with a timing resolution of 0.4 ns for event-per-event ion counting.
@@
Example spectra of the Doppler-corrected resonances observed when probing the transition for the even-even () isotopes are presented in Fig. 3, revealing the isotopic shift which contains the information of the changes in the nuclear charge radii .
@@
Throughout the experiment, the ion beam acceleration voltage was measured, allowing accurate Doppler-correction of the spectra. The wavelength of the scanned step was regularly measured using a WSU-2 HighFinesse wavemeter, which simultaneously recorded a reference wavelength from a stabilized laser diode (Toptica DLC DL PRO 780) locked to the Doppler-free hyperfine transition of 87Rb (TEM CoSy) for correcting drifts of the wavemeter’s interferometer. Furthermore, data taking on the isotopes of interest was interleaved throughout the experiment campaign with measurements of the selected reference isotope 112Sn to ensure any additional drifts were taken into account when determining the isotope shifts. {NoHyper}
| CRIS | COLLAPS | ||||||
| (MHz) | (MHz) | (MHz) | (MHz) | ||||
| 104 | 0 | ||||||
| 105 | + | ||||||
| 106 | 0 | ||||||
| 107 | + | ||||||
| 108 | 0 | ||||||
| 109 | + | ||||||
| 110 | 0 | ||||||
| 111 | + | ||||||
| 112 | 0 | 0 | 0 | ||||
| 113 | + | ||||||
| 114 | 0 | ||||||
| 115 | + | ||||||
| 116 | 0 | ||||||
| 117 | + | ||||||
| 118 | 0 | ||||||
| 119 | + | ||||||
| 120 | 0 | ||||||
| 121 | + | ||||||
| 122 | 0 | ||||||
| 123 | - | ||||||
| 124 | |||||||
| 125 | - | ||||||
| 126 | 0 | ||||||
| 127 | - | ||||||
| 128 | 0 | ||||||
| 129 | + | ||||||
| 130 | |||||||
| 131 | + | ||||||
| 132 | |||||||
| 133 | - | ||||||
| 134 | 0 | ||||||
@@
The (284 nm) transition in tin was predominantly used throughout the experiment, allowing spectroscopy of 21 tin isotopes ranging from 104Sn to 124Sn. Complementary measurements were performed using the (281 nm) transition down to 106Sn. During the spectroscopy of 114,116Sn using the transition, the laser power of the first step was increased by a factor five for enhancing ionization efficiency while increasing the linewidth from 230 MHz to 350 MHz FWHM. Reference measurements of both 112Sn and 120Sn were performed under these identical conditions. Resonances were fitted using a Lorentz profile with free linewidth, since no Gaussian contribution was observed, with an average linewidth of 230 MHz for all isotopes except those acquired using a higher laser power. As an example, the fitted spectra extracted using the transition of the even-even isotopes are presented in Fig. 3.
@@
The determined isotope shifts relative to 124Sn are provided in Table 2 for both transitions. In order to obtain from the isotope shifts measured at CRIS relative to 112Sn, the following procedure was used. First, the isotope shifts relative to 124Sn were calculated and the statistical uncertainties of the isotope of interest and that of the new reference 124Sn were added in quadrature. To determine the total uncertainty, we combined the statistical uncertainty with the systematic uncertainty specific to each isotope. In this particular case, the systematic uncertainty was not derived from a potential voltage offset. Instead, it stemmed from laser drifts that could occur during the measurement of different isotopes, whether in positive or negative directions. For the isotope 112Sn, we used the systematic uncertainty associated with 124Sn. This total uncertainty was then used to perform the King plot for the individual transitions. For the COLLAPS data, only the statistical uncertainty was taken into account, since the systematic uncertainty is dominated by the scaling uncertainty of the voltage divider that transfers into a mass-shift like behavior in the isotopic chain and is, thus, captured by the King-plot procedure.
@@
The charge radii of the stable isotopes for the King plots are obtained based on the data provided in the book of Fricke and Heilig [50-Sn]: The Barrett radii determined from muonic atom spectroscopy were divided by the isotope-dependent factors () obtained from elastic electron scattering to obtain the mean-square charge radii. From these, the of the stable even isotopes were obtained. We note that two sets of Barrett radii are listed in [50-Sn], which have been treated identically and the changes in the mean-square charge radii were finally averaged to arrive at a single set. These are listed in Table 2 and used in the four King plots.
The final value for as listed in the main manuscript in Table I is then obtained as the weighted mean of the King-plot results of the four transitions. The final uncertainty has been chosen as the smallest uncertainty that was obtained from one specific transition, but no further reduction was considered due to the strong correlations via the respective field shift uncertainty. It was found that the size of this uncertainty covers approximately the observed range of variation of the results in the four transitions.
The mass shift and field shift factors obtained in the King plots, which are required to obtain the values in the respective transitions, are summarized in Table LABEL:table:atomic_factors. The differential mean-square charge radii and their uncertainties for each transition individually can be obtained using the values in Table 2 and Table LABEL:table:atomic_factors, according to with the mass factor . In the evaluation of the uncertainty on , the optimized ensures that no correlations between and exist in the error propagation.